Fc=mv2/r
Fc is centripetal force, m is mass, v is speed and r is radius


6. What will happen to the size of the centripetal force if the mass (m) is doubled?


1)No change
2)The centripetal force doubles
3)The centripetal force increases by a factor of 4
4)The centripetal force is halved
1
Expert's answer
2019-10-10T09:56:34-0400

The centripetal force


"F=\frac{mv^2}{r}"

If the mass is doubled then


"F'=\frac{2mv^2}{r}=2\times\frac{mv^2}{r}=2F"

Answer: (2) The centripetal force doubles.